Find the general solution of the differential equations taking into account the initial conditions using the parameter variation method:

yiv + 2y + y = 3t + 4 ; y(0) = y(0) = 0 et y(0) = y(0) = 1
Answer #1

Date: Saint y + 24 +y = 36 = 3t +4 the Auxilery ogor at the defp eq? is given hy m +2m tlzo (m²+1) → m2+1 zo -] AN ma m = >Daic: 7.c3fSina dut) -62 Sint 3 I gott) 3t ty Hence the general Sdn given by yh- (G that) Cast + ((3+ (4) Sint + 3t +4 Ho)=0ك Date: / / =)) -3-(- ) تو) L - 3( - 22 2-( (2 7 5 =72 ا 3 hence the general soin .ک y Lt) + ا ره. پرت + 4 - + (.% - 3+) ne +